OpenStudy (anonymous):

Solve: x2 + 17x = -30 {15, 2} {-15, -2} {5, 6} {-5, 6}

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

(-15,-2)

OpenStudy (anonymous):

x2+17x=-30 -15,-2 is correct...

OpenStudy (anonymous):

x2+17x+30=0 (x+15)(x+2)=0 x= -15 and x= -2

OpenStudy (anonymous):

duh

OpenStudy (anonymous):

Factor completely: 14x2 + 47x + 15 (14x + 3)(x + 5) (14x + 5)(x + 3) (2x + 3)(7x + 5) (2x + 5)(7x + 3)